OpenStudy (anonymous):

Will someone please help me I posted this like and hour ago and closed it because no one replied :( help please the halfway point of a tunnel through a mountain is 3/2 miles from either end of the tunnel. the mountain is 660 feet (1/8 mile) high. Find the slope of. the side of the mountain

OpenStudy (zzr0ck3r):

For this we use the Pythagorean theorem. We have something like this.  we use \((\text{slope})^2=(\frac{3}{2})^2+(\frac{1}{8})^2\) solving for slope we get \(\text{slope}=\sqrt{(\frac{3}{2})^2+(\frac{1}{8})^2}\)